# A proof of the reverse isoperimetric inequality using a geometric–analytic approach

### v2

@@ II.2 Effective Functional and Its Variation
$\begin{split}&\Phi_{TT}\;=\;-\,\frac{C_{\Omega,V}}{T}\;<\;0,\qquad\Phi_{\Omega\Omega}\;=\;-\,\chi_{T,V}\;<\;0,\\ &\Delta\;\equiv\;\det\!\begin{pmatrix}\Phi_{TT}&\Phi_{T\Omega}\\ \Phi_{T\Omega}&\Phi_{\Omega\Omega}\end{pmatrix}\;>\;0,\end{split}$ (45)
where $C_{\Omega,V}$ is the specific heat at fixed $(\Omega,V)$ and $\chi_{T,V}$ is the isothermal moment of inertia at fixed $(T,V)$; both are positive on the stable branch.
@@
the map $(T,\Omega)\mapsto(S,J)$ is invertible, since its Jacobian is $\Delta$. One can then express the curvature of $S$ as a function of $J$ at fixed $V$ entirely in terms of $\Phi$:
$\left(\frac{\partial^{2}S}{\partial J^{2}}\right)_{V}\;=\;-\frac{\Delta}{\Phi_{TT}\,\Phi_{\Omega\Omega}}\;<\;0.$ (46)
This exact identity shows that $S(V,J)$ is strictly concave in $J$ along the connected thermodynamically stable branch. Combining this global concavity with the fact that $J=0$ is a local maximum of entropy implies
$S(V,J)\;\leq\;S(V,0)\,,\quad\text{with strict inequality for }J\neq 0\,$, (47)
demonstrating that rotation lowers the entropy at fixed thermodynamic volume and $V$ that the Schwarzschild–AdS solution ($J=0$) is the global maximizer along the Kerr–AdS family. The above inequality holds for arbitrary $J$ along the entire connected thermodynamically stable branch.
@@
The proof presented here establishes the Reverse Isoperimetric Inequality for the class of stationary, asymptotically AdS black holes in $D\geq 4$ Einstein gravity, with or without Maxwell fields at fixed charge $Q$, under the following conditions: (i) the horizon is compact, connected, and of spherical topology; (ii) variations are performed in the fixed–$V$ (and fixed–$Q$) ensemble; and (iii) the analysis is restricted to Einstein gravity, so that the thermodynamic quantities are governed by the standard Einstein–Hilbert action and Gibbons–Hawking–York boundary terms.
@@
III Conclusion and Discussion
In this work, we have established a rigorous geometric–analytic proof of the reverse isoperimetric inequality (RII) for black holes in Einstein gravity for $D\geq 4$ AdS space. On the one hand, the geometric argument shows using gravitational focusing and Sherif–Dunsby rigidity theorem that a round Euclidean horizon $S^{D-2}$ enclosing a fixed thermodynamic volume is a maximally entropic surface. On the other hand, the analytic approach demonstrates this using the second variation of entropy; volume-preserving ($\ell\geq 2$) perturbations of the round sphere yield $\delta^{2}\mathcal{A}<0$ in AdS space, confirming that the round AdS-Schwarzschild horizon is a local maximum of area (entropy) under volume constraints. We then extended our analysis to include a Kerr-AdS black hole and confirmed that it is not an entropy maximizer at a fixed thermodynamic volume.
@@ III Conclusion and Discussion
  • Effects of Modified Gravity on the RII: In theories beyond Einstein gravity, such as $f(R)$, Gauss–Bonnet or more general Lovelock gravities, and scalar–tensor models, the gravitational field equations acquire extra curvature‐dependent or scalar‐coupling terms which modify both the Raychaudhuri focusing condition and the form of the Euclidean action functional. In particular, the sheet‐expansion scalar $\theta$ no longer obeys the simple $\theta<0$ condition under volume‐preserving deformations. Analytically, the second variation of $I_{\rm slice}$ acquires extra terms $\delta^{2}I_{\rm higher\text{-}curv}$ coming from variation of the $f(R)$ or Gauss–Bonnet invariants, altering the stability criterion of the extremal hypersurface. A systematic study of these corrections would therefore be required in alternative theories of gravity.
  • Quantum Corrections: Incorporating higher-curvature corrections or quantum effects (e.g., via one-loop determinants or entanglement entropy corrections) may modify the geometric rigidity or the second variation functional, potentially leading to refined “quantum RII” bounds.
  • Holographic Perspectives: Given the AdS/CFT correspondence, it would be interesting to interpret our RII proof in the dual field theory, perhaps relating maximal horizon entropy at fixed volume to extremal entanglement or energy constraints in the boundary CFT.
  • Beyond Asymptotic AdS: Extending the analysis to asymptotically flat or more exotic asymptotics (e.g., Lifshitz, hyperscaling violation) might reveal whether the reverse isoperimetric phenomenon is unique to constant-$\Lambda$ backgrounds or has broader applicability. Moreover, the RII has been advocated to hold for dS black holes in [21]. So, it is interesting to see if the proof can be extended to this case.
  • Violation in the case of superentropic black holes: It is known that superentropic black holes violate RII. As part of the proof we presented, this can be traced to their non-compact hypersurface, while the proof requires a compact hypersurface. Nevertheless, a general proof explicitly for non-compact hypersurfaces is an interesting future work.
@@ Appendix A Conformal Rigidity and Sphericity of Compact Hypersurfaces
  • Compactness and positive scalar curvature. Each hypersurface $\mathcal{T}\simeq S^{3}$ is compact and, by the Yamabe theorem, admits a conformal representative of constant scalar curvature. Since $\mathcal{T}$ lies in an AdS (Einstein) background with negative cosmological constant, the Gauss–Codazzi relation ensures $R(h)>0$ on every spatial slice. Moreover, the event–horizon topology theorem [22] implies that stationary black–hole horizons are of positive Yamabe type, confirming the required curvature sign.
  • Non–Einstein character. While the ambient spacetime $(\mathcal{M},g)$ satisfies $R_{ab}=\Lambda g_{ab}$, the deformed slice $\mathcal{T}$ need not: its intrinsic Ricci tensor is not proportional to $h_{ab}$, i.e. it is genuinely non–Einstein. This precisely corresponds to the condition $\alpha\neq\beta$ in the theorem.
  • Scalar–curvature–preserving conformal deformation. Within the conformal class $h_{ab}]$, one may choose a representative $\tilde{h}_{ab}=e^{2\varphi}h_{ab}$ satisfying
@@
    This condition freezes the intrinsic scalar curvature, ensuring that variations probe only the extrinsic shape of the hypersurface rather than its internal curvature profile.
  • Negative conformal factor from gravitational focusing. In AdS, $R_{ab}e^{a}e^{b}=\Lambda<0$, and the Raychaudhuri equation for the sheet expansion $\theta=\delta_{\mu}e^{\mu}$ implies $\theta<0$ everywhere on $\mathcal{T}$. In the 1+1+21{+}1{+}2 formalism this is equivalent to a negative conformal factor $\varphi<0$ in the transformation $\tilde{h}_{ab}=e^{2\varphi}h_{ab}$. Thus, the “proper” conformal rescaling required by the theorem is precisely realized by gravitational focusing of the normal congruence.

### v3

@@ III.2 Concavity of $S$ in $J$ at fixed thermodynamic volume $\mathcal{V}$
$\begin{split}&\Phi_{TT}\;=\;-\,\frac{C_{\Omega,\mathcal{V}}}{T}\;<\;0,\qquad\Phi_{\Omega\Omega}\;=\;-\,\chi_{T,\mathcal{V}}\;<\;0,\\ &\Delta\;\equiv\;\det\!\begin{pmatrix}\Phi_{TT}&\Phi_{T\Omega}\\ \Phi_{T\Omega}&\Phi_{\Omega\Omega}\end{pmatrix}\;>\;0,\end{split}$ (56)
where $C_{\Omega,\mathcal{V}}$ is the specific heat at fixed $(\Omega,\mathcal{V})$ and $\chi_{T,\mathcal{V}}$ is the isothermal moment of inertia at fixed $(T,\mathcal{V})$; both are positive on the stable branch.
@@
the map $(T,\Omega)\mapsto(S,J)$ is invertible, since its Jacobian is $\Delta$. One can then express the curvature of $S$ as a function of $J$ at fixed $\mathcal{V}$ entirely in terms of $\Phi$:
$\left(\frac{\partial^{2}S}{\partial J^{2}}\right)_{\mathcal{V}}\;=\;-\frac{\Delta}{\Phi_{TT}\,\Phi_{\Omega\Omega}}\;<\;0.$ (57)
This exact identity shows that $S(\mathcal{V},J)$ is strictly concave in $J$ along the connected thermodynamically stable branch. Combining this global concavity with the fact that $J=0$ is a local maximum of entropy implies [^7] The Reverse Isoperimetric Inequality is a $\mathcal{V}$-only statement: $S(\mathcal{V})\leq S_{\mathrm{Schw}}(\mathcal{V})$. In the thermodynamic subsection, we temporarily resolve the implicit $J$-dependence by writing $S(\mathcal{V},J)$ (equivalently $S_{\mathcal{V}}(J)$) for the entropy of the equilibrium state with fixed thermodynamic volume $\mathcal{V}$ and angular momentum $J$. Here, $J$ merely labels deformations within the fixed-$\mathcal{V}$ Kerr–AdS branch; it is not an independent argument of the RII. Accordingly, statements such as strict concavity $d^{2}S_{\mathcal{V}}/dJ^{2}<0$ (implying $S_{\mathcal{V}}(J)\leq S_{\mathcal{V}}(0)$) serve as consistency checks and do not alter the $\mathcal{V}$-only content of the inequality.
[…]
demonstrating that rotation lowers the entropy at fixed thermodynamic volume $\mathcal{V}$ and that the Schwarzschild–AdS solution ($J=0$) is the global maximizer along the Kerr–AdS family. The above inequality holds for arbitrary $J$ along the entire connected thermodynamically stable branch.
@@
[…]
@@
IV Conclusion and Discussion
In this work, we have established a geometric–analytic proof of the reverse isoperimetric inequality (RII) for black holes in Einstein gravity in AdS spacetime for $D\geq 4$. On the one hand, the geometric argument shows, using gravitational focusing together with the Sherif–Dunsby rigidity theorem, that the round horizon $S^{D-2}$ enclosing a fixed thermodynamic volume maximizes the entropy within the class of pure shape deformations considered. On the other hand, the analytic approach demonstrates this using the second variation of entropy: nontrivial volume-preserving ($\ell\geq 2$) perturbations of the round sphere yield $\delta^{2}\mathcal{A}<0$ in AdS space, confirming that the AdS–Schwarzschild horizon is a local maximum of area (entropy) under volume constraints, with the geometric argument providing the corresponding global upgrade within this deformation class. We then extend the analysis to Kerr–AdS black holes and show that they do not maximize the entropy at fixed thermodynamic volume.
@@ IV Conclusion and Discussion
Effects of Modified Gravity on the RII.
In theories beyond Einstein gravity, such as $f(R)$, Gauss–Bonnet or more general Lovelock gravities, and scalar–tensor models, the gravitational field equations acquire extra curvature‐dependent or scalar‐coupling terms which modify both the Raychaudhuri focusing condition and the form of the Euclidean action functional. In particular, the sheet‐expansion scalar $\theta$ no longer obeys the simple $\theta<0$ condition under volume‐preserving deformations. Analytically, the second variation of $I_{\rm slice}$ acquires extra terms $\delta^{2}I_{\rm higher\text{-}curv}$ coming from variation of the $f(R)$ or Gauss–Bonnet invariants, altering the stability criterion of the extremal hypersurface. A systematic study of these corrections would therefore be required in alternative theories of gravity.
Quantum Corrections.
Incorporating higher-curvature corrections or quantum effects (e.g., via one-loop determinants or entanglement entropy corrections) may modify the geometric rigidity or the second variation functional, potentially leading to refined “quantum RII” bounds.
Holographic Perspectives.
Given the AdS/CFT correspondence, it would be interesting to interpret our RII proof in the dual field theory, perhaps relating maximal horizon entropy at fixed volume to extremal entanglement or energy constraints in the boundary CFT.
Beyond Asymptotic AdS.
Extending the analysis to asymptotically flat or more exotic asymptotics (e.g., Lifshitz, hyperscaling violation) might reveal whether the reverse isoperimetric phenomenon is unique to constant-$\Lambda$ backgrounds or has broader applicability. Moreover, the RII has been advocated to hold for dS black holes in [6]. So, it is interesting to see if the proof can be extended to this case.
Violation in the case of superentropic black holes.
It is known that superentropic black holes violate RII. As part of the proof we presented, this can be traced to their non-compact hypersurface, while the proof requires a compact hypersurface. Nevertheless, a general proof explicitly for non-compact hypersurfaces is an interesting future work.
@@ Appendix A Conformal Rigidity and Sphericity of Compact Hypersurfaces
Compactness and positive scalar curvature.
Each hypersurface $\mathcal{T}\simeq S^{3}$ is compact and, by the Yamabe theorem, admits a conformal representative of constant scalar curvature. Since $\mathcal{T}$ lies in an AdS (Einstein) background with negative cosmological constant, the Gauss–Codazzi relation ensures $R(h)>0$ on every spatial slice. Moreover, the event–horizon topology theorem [12] implies that stationary black–hole horizons are of positive Yamabe type, confirming the required curvature sign.
Non–Einstein character.
While the ambient spacetime $(\mathcal{M},g)$ satisfies $R_{ab}=\Lambda g_{ab}$, the deformed slice $\mathcal{T}$ need not: its intrinsic Ricci tensor is not proportional to $h_{ab}$, i.e. it is genuinely non–Einstein. This precisely corresponds to the condition $\alpha\neq\beta$ in the theorem.
Scalar–curvature–preserving conformal deformation.
Within the conformal class $h_{ab}]$, one may choose a representative $\tilde{h}_{ab}=e^{2\varphi}h_{ab}$ satisfying
@@
This condition freezes the intrinsic scalar curvature, ensuring that variations probe only the extrinsic shape of the hypersurface rather than its internal curvature profile. This allows the geometric argument to probe whether a non-round shape can maximize entropy at fixed thermodynamic volume without introducing additional curvature effects.
Negative conformal factor from gravitational focusing.
In an Einstein background with negative cosmological constant, $R_{ab}=\Lambda g_{ab}$ with $\Lambda<0$. The Raychaudhuri equation for the sheet expansion $\theta=\delta_{\mu}e^{\mu}$ implies gravitational focusing of the spacelike sheet congruence generated by $e^{a}$, driving the congruence toward a contracting regime along the flow. We therefore evaluate the rigidity theorem once the congruence enters the regime
$\theta<0,$ (62)
in which the 1+1+21{+}1{+}2 formalism identifies the geometry with a contracting conformal rescaling of the sheet metric. In particular, the intrinsic metric on $\mathcal{T}$ admits a conformal representative
$\tilde{h}_{ab}=e^{2\varphi}h_{ab},\qquad\varphi<0,$ (63)
which preserves the scalar curvature. Thus the “proper” conformal transformation required by the Sherif–Dunsby theorem arises naturally from the gravitational focusing of the normal congruence.
@@ Appendix C Extrinsic and mean curvature of a 2–sphere embedded in Euclidean AdS4
Appendix D Rotating generalization of Dolan’s identification $I_{E}=\beta G$
In this Appendix we generalize the standard identification of the Euclidean on–shell gravitational action with a thermodynamic potential (as emphasized for the static case by Dolan [7]) to the rotating grand–canonical ensemble. The key output is the rotating analogue of the familiar relation $I_{E}=\beta F$:
$I_{E}=\beta\,G(T,\Omega,P),\qquad G\equiv M-TS-\Omega J,$ (80)
so that, in extended thermodynamics,
$\left(\frac{\partial I_{E}}{\partial P}\right)_{T,\Omega}=\beta\,V,\qquad V=\left(\frac{\partial G}{\partial P}\right)_{T,\Omega}.$ (81)
This shows that the volume extracted from the Euclidean action is the standard thermodynamic volume appearing in the Lorentzian first law/Smarr relation (and hence in the Komar/Iyer–Wald formulation used in black-hole chemistry).
Rotating grand–canonical boundary data
We consider stationary, asymptotically AdS saddles in Euclidean signature, with boundary data fixed at infinity
$(\beta,\Omega,\Lambda)\qquad\text{equivalently}\qquad(T,\Omega,P),\qquad P\equiv-\frac{\Lambda}{8\pi G}.$ (82)
The Euclidean time circle is generated by the Killing vector
$\xi\;=\;\partial_{\tau}+\Omega\,\partial_{\phi},$ (83)
where $\Omega$ is the angular velocity defined with respect to a non–rotating frame at infinity (the standard convention in extended black-hole thermodynamics). To ensure the metric remains real and positive-definite (Riemannian) in the Euclidean section, we employ the standard analytic continuation $t\to-i\tau$ and $J\to iJ$ (implying $a\to ia$ and $\Omega\to i\Omega_{E}$), as detailed in [7].
The corresponding grand–canonical partition function is
[…]
so semiclassically
$Z(\beta,\Omega,P)\sim e^{-I_{E}[g_{\rm cl}]},$ (85)
with $g_{\rm cl}$ the Euclidean saddle obeying the boundary conditions (82).
Euclidean action and the Hamiltonian generator $H_{\xi}=M-\Omega J$
We start from the regulated Euclidean Einstein action (including the Gibbons–Hawking–York term and the usual AdS counterterms)
$\begin{split}&I_{E}[g]=-\frac{1}{16\pi G}\int_{M}(R-2\Lambda)\sqrt{g}\,d^{4}x-\frac{1}{8\pi G}\int_{\partial M}K\sqrt{\gamma}\,d^{3}x\\ &\qquad+I_{\rm ct}[\gamma].\end{split}$ (86)
For a stationary saddle, the Hamiltonian (ADM) decomposition expresses the on–shell action as a pure boundary term plus a horizon contribution. The boundary term is the conserved charge associated with the Euclidean time generator $\xi$, i.e. the Hamiltonian $H_{\xi}$ that generates translations along (83). In the rotating grand–canonical ensemble this is
$H_{\xi}=M-\Omega J,$ (87)
because $\xi$ combines time translations with a rigid rotation at infinity.
Accordingly, the on–shell Euclidean action takes the universal form
$I_{E}=\beta\,H_{\xi}-S\;=\;\beta\,(M-\Omega J)-S.$ (88)
The entropy term $S$ arises from the horizon (bolt) contribution, exactly as in the static case: it is fixed by regularity of the Euclidean section and equals $A_{H}/(4G)$. Importantly, rotation changes the boundary generator from $M$ to $M-\Omega J$ but does not change the structure of the horizon term.
Equation (88) is the rotating generalization of the standard static identification discussed by Dolan [7].
Identification with Gibbs free energy
Define the Gibbs free energy in the rotating grand–canonical ensemble:
$G(T,\Omega,P)\equiv M-TS-\Omega J.$ (89)
Using $\beta=1/T$, the result (88) immediately gives
$I_{E}=\beta\,G,$ (90)
which establishes (80).
Thermodynamic consequence: $(\partial I_{E}/\partial P)_{T,\Omega}=\beta V$
From the extended first law
$dM=T\,dS+\Omega\,dJ+V\,dP,$ (91)
we obtain
$dG=d(M-TS-\Omega J)=-S\,dT-J\,d\Omega+V\,dP.$ (92)
Multiplying by $\beta$ and using $I_{E}=\beta G$ yields
$dI_{E}=\beta\,dG=\beta\,(V\,dP-S\,dT-J\,d\Omega),$ (93)
and therefore at fixed $(T,\Omega)$,
$\left(\frac{\partial I_{E}}{\partial P}\right)_{T,\Omega}=\beta\,V.$ (94)
This is the precise sense in which the “volume extracted from the Euclidean action” coincides with the standard thermodynamic volume appearing in the Lorentzian first law and Smarr relation. In particular, for Kerr–AdS black holes this $V$ is the same thermodynamic (Komar/Iyer–Wald) volume employed in black-hole chemistry and in the formulation of the reverse isoperimetric inequality.
Explicit Euclidean derivation of the Kerr–AdS thermodynamic volume
In this subsection we explicitly demonstrate, for rotating black holes, that the pressure derivative of the Euclidean on-shell action reproduces the standard Kerr–AdS thermodynamic volume, without assuming a priori the identification $I_{E}=\beta G$. This provides a direct Euclidean confirmation of the Lorentzian extended thermodynamics.
We work in four dimensions and follow the Euclidean action computation of Cheng et al. [3]. The regulated Euclidean action for Kerr–AdS, evaluated on a smooth saddle (with no conical defect), is
$\begin{split}I_{E}(r_{+},a,L;\beta)&=\frac{\beta\,r_{+}}{2G\,\Xi}\left(1+\frac{r_{+}^{2}}{L^{2}}\right)-\frac{\pi(r_{+}^{2}+a^{2})}{G\,\Xi},\\ \Xi&\equiv 1-\frac{a^{2}}{L^{2}},\end{split}$ (95)
where $r_{+}$ is the horizon radius, $a$ the rotation parameter, and $L$ the AdS length. The inverse temperature $\beta$ is fixed by regularity of the Euclidean section.
The rotating grand-canonical ensemble is defined by fixing the boundary data $(\beta,\Omega)$, where the angular velocity measured with respect to a non-rotating frame at infinity is
$\Omega=\frac{a}{r_{+}^{2}+a^{2}}\left(1+\frac{r_{+}^{2}}{L^{2}}\right).$ (96)
The Hawking temperature is
$\begin{split}T_{H}=\frac{r_{+}}{4\pi(r_{+}^{2}+a^{2})}\left(1+\frac{a^{2}}{L^{2}}+\frac{3r_{+}^{2}}{L^{2}}-\frac{a^{2}}{r_{+}^{2}}\right),\end{split}$ (97)
so that $\beta=T_{H}^{-1}$.
We now vary the cosmological constant, or equivalently the AdS length $L$, while holding the ensemble data fixed:
$\delta\beta=0,\qquad\delta\Omega=0.$ (98)
These conditions induce variations of $r_{+}$ and $a$ when $L$ is varied. At linear order the constraints read
$\begin{split}\partial_{r_{+}}T_{H}\,dr_{+}+\partial_{a}T_{H}\,da+\partial_{L}T_{H}\,dL&=0,\\ \partial_{r_{+}}\Omega\,dr_{+}+\partial_{a}\Omega\,da+\partial_{L}\Omega\,dL&=0.\end{split}$ (99)
For later use, the required partial derivatives are
$\begin{split}&\partial_{r_{+}}\Omega=-\frac{2ar_{+}\Xi}{(r_{+}^{2}+a^{2})^{2}},\\ &\partial_{a}\Omega=\frac{(1+r_{+}^{2}/L^{2})(r_{+}^{2}-a^{2})}{(r_{+}^{2}+a^{2})^{2}},\qquad\partial_{L}\Omega=-\frac{2ar_{+}^{2}}{L^{3}(r_{+}^{2}+a^{2})},\end{split}$ (100)
and
$\partial_{a}T_{H}=-\frac{ar_{+}(L^{2}+r_{+}^{2})}{\pi L^{2}(r_{+}^{2}+a^{2})^{2}},\qquad\partial_{L}T_{H}=-\frac{r_{+}(a^{2}+3r_{+}^{2})}{2\pi L^{3}(r_{+}^{2}+a^{2})},$ (101)
with $\partial_{r_{+}}T_{H}$ determined straightforwardly.
Solving (99) yields $dr_{+}/dL$ and $da/dL$, which are then inserted into the total variation of the Euclidean action at fixed $\beta$,
$dI_{E}=\partial_{r_{+}}I_{E}\,dr_{+}+\partial_{a}I_{E}\,da+\partial_{L}I_{E}\,dL.$ (102)
The partial derivatives of $I_{E}$ at fixed $\beta$ are
$\begin{split}&\partial_{r_{+}}I_{E}=\frac{\beta(L^{2}+3r_{+}^{2})-4\pi L^{2}r_{+}}{2GL^{2}\Xi},\\ &\partial_{a}I_{E}=\frac{a(L^{2}+r_{+}^{2})(\beta r_{+}-2\pi L^{2})}{GL^{4}\Xi^{2}},\end{split}$ (103)
$\partial_{L}I_{E}=-\frac{(r_{+}^{2}+a^{2})(\beta r_{+}-2\pi a^{2})}{GL^{3}\Xi^{2}}.$ (104)
We now express the variation in terms of the thermodynamic pressure, defined in four dimensions by
$P=\frac{3}{8\pi GL^{2}},\qquad\frac{dL}{dP}=-\frac{4\pi GL^{3}}{3}.$ (105)
Combining these results and finally imposing the smoothness condition $\beta=T_{H}^{-1}(r_{+},a,L)$, one finds
$\left(\frac{\partial I_{E}}{\partial P}\right)_{\beta,\Omega}=\beta\,V,$ (106)
where
$\begin{split}V=\frac{4\pi r_{+}(r_{+}^{2}+a^{2})}{3\,\Xi}\left[1+\frac{a^{2}}{2r_{+}^{2}}\frac{1+r_{+}^{2}/L^{2}}{\Xi}\right].\end{split}$ (107)
The expression (107) coincides exactly with the Kerr–AdS thermodynamic volume obtained from the Lorentzian extended first law and Smarr relation [5]. Thus, in the rotating grand-canonical ensemble, the Euclidean on-shell action reproduces the correct thermodynamic volume directly through its pressure derivative, without any independent assumption regarding its interpretation as a Gibbs free energy.